OpenStudy (anonymous):

The sum of x and y numbers is 20. The difference of the two numbers is 16. If the numbers are x and y, which system of linear equations represents the problem?

OpenStudy (stamp):

The sum of x and y numbers is 20. sum means you add. so les think The sum of x and y numbers is 20. x + y thats what the sum means x + y IS 20 is?? what does is mean x + y = 20

OpenStudy (stamp):

The difference of the two numbers is 16. difference means you subtract so The difference of the two numbers is 16. the two numbers are x and y do you remember? so The difference of the two numbers is 16. x - y thats what the difference is x -y is 16 do you remember what is means is means it IS x - y = 16

OpenStudy (stamp):

ok so x + y = 20 but x - y = 16 so X + Y = 20 X - Y = 16

OpenStudy (anonymous):

well that was easy for u it took me to try's I'm not really good at math thank you

OpenStudy (anonymous):

I mean two ( 2 )